OpenStudy (anonymous):

Find the probability for the experiment of drawing two marbles (without replacement) from a bag containing three green, four yellow, and five red marbles such that the marbles are different colors.

OpenStudy (anonymous):

I got 38/121, but that isn't a choice on my multiple choice question. The answer has to be 47/132, 47/30, 47/66, 1/6, or 2/3

OpenStudy (anonymous):

and it can't be 47/30, because it has to be 0<x<1

OpenStudy (amistre64):

3 colors g g g y g r y y y g y r r r r y r g 9 ways to draw out 2 marbles 6 ways in which they are different collors 6/9 = 2/3, if i see it right

OpenStudy (anonymous):

your guess is better than mine

OpenStudy (amistre64):

but you might have to prob up the picks and some fancier stuff :)

OpenStudy (anonymous):

yeah, i have no idea

OpenStudy (anonymous):

it was 47/66 as the answer

OpenStudy (amistre64):

3g, 4y, 5r ; 12 total g g 3*2 / 132 g y 3*4 / 132 g r 3*5 / 132 y y 4*3 / 132 y g 4*3/ 132 y r 4*5/ 132 r r 5*5/132 r y 5*4/132 r g 5*3/132 P(g,y or gr or ...) 12 +15 +12 +20 +20 +15 = 24 +30+40 = 24+70 = 94 94/132 = 47/66 is possible as well